OpenStudy (anonymous):

An ice cream cone is left sitting in the hot sun. Sarah notices that the ice cream melts and loses half of its volume every 5 minutes. If the starting volume was 125 mL, determine a function for the volume, with respect to the amount of time left out in the sun.

OpenStudy (anonymous):

Hi all. I have no idea how to do this.

imqwerty (imqwerty):

hey robert

OpenStudy (anonymous):

Hey qwerty

OpenStudy (anonymous):

Long time no see

imqwerty (imqwerty):

hahaha :D welcome back

OpenStudy (anonymous):

Thanks bro

OpenStudy (anonymous):

I'm traveling through Venezuela now

imqwerty (imqwerty):

cool :D ok so the volume is expanding it will be given like this-\[V=V_o(1+\gamma \Delta T)\]here γ=coefficient of 3D expansion V_o is the initial volume V=final volume

OpenStudy (anonymous):

Oh ok.

OpenStudy (anonymous):

Let me see if I understand the equation

imqwerty (imqwerty):

i am not so sure i am also currently studying this topic :) but i think its ok I have to leave now i have to go somewhere

OpenStudy (baru):

i think its exponential decay

OpenStudy (anonymous):

It is an exponential decay

OpenStudy (baru):

5 min is the half life

OpenStudy (anonymous):

Correct. But it's been a long while since I did exponential decay

OpenStudy (anonymous):

I think Qwerty's solution seems solid but that equation needs to be substituted with values I believe.

imqwerty (imqwerty):

is this related to solid dynamics?

OpenStudy (anonymous):

It's related to advanced functions.

OpenStudy (baru):

i first came across it in context of radioactive elements

OpenStudy (anonymous):

And I'm sure physics explains better but this is some math jam

OpenStudy (anonymous):

How may I express every five minutes part in exponential decay equation? My skills are decayed.

imqwerty (imqwerty):

wait a sec ima attach some of my notes related to this

OpenStudy (anonymous):

Oh Thanks bro

ganeshie8 (ganeshie8):

Try : \[V(t) = 125*(1/2)^{t/5}\]

OpenStudy (anonymous):

That sums up everything....

ganeshie8 (ganeshie8):

when t=5, we get : V(5) = 125*(1/2)^1 = 125/2 so the volume does get halved and we're good

OpenStudy (anonymous):

Wow @ganeshie8 how can you think of math like a language?

ganeshie8 (ganeshie8):

its easy to guess that function by trial and error haha

OpenStudy (baru):

@ganeshie8 thats very interesting!! why is it that radioactive decay is modelled by an exponential function?

OpenStudy (anonymous):

Ahh I see. I really need to get the differential equation going.

ganeshie8 (ganeshie8):

They are same thing. We could also solve a differential equation and get the same answer. The rate of change of volume is proportional to the existing volume. It is -1/2V per 5 minutes, which is same as -1/10V per minute : \[\dfrac{dV}{dt} = -\dfrac{V}{10}\]

OpenStudy (baru):

they are the samething? both the exponential funct and yours solve the differential equation?

OpenStudy (anonymous):

I think there are many ways of going about this problem.

OpenStudy (anonymous):

The one Ganesha introduced me at the beginning is more elementary

ganeshie8 (ganeshie8):

\[2^r = e^{r*\ln 2}\]

OpenStudy (anonymous):

And one with differential equation is more advanced and uses mathematical logics.

OpenStudy (anonymous):

Scott wants to calculate the distance from his house to each of his friends' houses. If he drives at 50 km/h, find a function for the distance, with respect to the number of hours it takes to travel.

OpenStudy (anonymous):

I think this can be expressed using differential equations too.

OpenStudy (anonymous):

F(x)=50km/h* x(h)

OpenStudy (anonymous):

Where x is the over all distance and oopps, the y being the hours taken

OpenStudy (anonymous):

But let's see how we can express this in differential equation.

ganeshie8 (ganeshie8):

that looks good, but we normally use "t" for time..

ganeshie8 (ganeshie8):

d(t) = 50km/h* t

OpenStudy (anonymous):

ah ok.. Sorry for my rusty math.... what a shame.

ganeshie8 (ganeshie8):

solving below differential equation gives the same function \[\dfrac{ds}{dt} = 50\]

OpenStudy (anonymous):

Ah ok that rings my bell.

OpenStudy (anonymous):

Thanks so much Ganesha!

OpenStudy (anonymous):

Math is wonderful thing you know

OpenStudy (baru):

\[e^{-kt}=(1/2)^{t/5}\] ??

ganeshie8 (ganeshie8):

Yes, \[k= \ln(1/2)/5\]

OpenStudy (baru):

phew!! so all is ok with the world thanks!!!

ganeshie8 (ganeshie8):

haha my previous reply has a mistake https://s3.amazonaws.com/upload.screenshot.co/8f51321707

ganeshie8 (ganeshie8):

v/10 is not a continuous change so we need to massage this expression further while setting up the de

OpenStudy (baru):

ohh..you mean you've got the rate constant wrong k=1/10 wrong?

ganeshie8 (ganeshie8):

yes, \(V(t) = 125e^{-kt} ~~\) we may use \(V(5) = 125/2\) to solve \(k\)

ganeshie8 (ganeshie8):

the differential equation should be simply \[\dfrac{dV}{dt} = -kV\] \(k\) is a constant which can be worked from the given info

OpenStudy (baru):

yes :) thanks again

ganeshie8 (ganeshie8):

np:)